# A Survey of World Models for Autonomous Driving

### v1

@@ Abstract
Recent breakthroughs in autonomous driving have revolutionized the way vehicles perceive and interact with their surroundings. In particular, world models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics. Such models unify perception, prediction, and planning, thereby enabling autonomous systems to make rapid, informed decisions under complex and often unpredictable conditions. Research trends span diverse areas, including 4D occupancy prediction and generative data synthesis, all of which bolster scene understanding and trajectory forecasting. Notably, recent works exploit large-scale pretraining and advanced self-supervised learning to scale up models’ capacity for rare-event simulation and real-time interaction. In addressing key challenges – ranging from domain adaptation and long-tail anomaly detection to multimodal fusion – these world models pave the way for more robust, reliable, and adaptable autonomous driving solutions. This survey systematically reviews the state of the art, categorizing techniques by their focus on future prediction, behavior planning, and the interaction between the two. We also identify potential directions for future research, emphasizing holistic integration, improved computational efficiency, and advanced simulation. Our comprehensive analysis underscores the transformative role of world models in driving next-generation autonomous systems toward safer and more equitable mobility.
@@ 1.1 Overview
Within this context, constructing robust and stable world models has emerged as a foundational element. The notion of a world model involves creating a high-fidelity representation of the driving environment – encompassing static structures (e.g., roads, buildings) and dynamic entities (e.g., vehicles, pedestrians) [3, 8]. A comprehensive world model continuously captures semantic and geometric information while updating these representations in real-time, thereby informing downstream tasks such as physical world prediction [12, 13]. Recent advances integrate multi-sensor data to refine these representations, such as generative approaches [14, 15] that simulate the physical world for training that unify heterogeneous sensor inputs into consistent top-down perspectives [16, 17].
@@ 1.2 Contributions
Guided by the principle that the world model is central to the understanding of dynamic scenes, this survey aims to provide a comprehensive, structured review of existing methodologies. We categorize state-of-the-art research into three key areas: Future Prediction of the Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [11, 25]; Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [12, 13]; Interaction Between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [26, 27, 18]. Specifically, we provide:
  • An In-Depth Analysis of Future Prediction Models: We discuss how Image-/BEV-/OG-/PC-based methods achieve geometric and semantic fidelity in dynamic scenes, including 4D occupancy forecasting and diffusion-based generation.
@@
§1§2§3§4§5§6§7§8§3.1§3.2§3.3§4
Figure 1: Structure of the overall review. The first row shows the structure of the paper. The second and third rows detail future prediction of the physical world, behavior planning for intelligent agents, and interaction between behavior planning and future prediction. The fourth row highlights various methodologies for training models in autonomous driving, covering both self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation.
@@ 1.3 Structure
A summary of the structure of this paper can be found in Fig. 1, which is presented as follows: Sec. 1 introduces the significance of world models in autonomous driving and outlines the societal and technical challenges they address. Sec. 2 provides a comprehensive background on the formulation and core tasks of world models in autonomous driving, specifically focusing on the future prediction of the physical world and behavior planning for intelligent agents. Sec. 3 details the taxonomy of methods: Sec. 3.1 delves into methods for future prediction of the physical world, discussing physical world evolution of dynamic objects and static entities. Sec. 3.2 discusses advanced behavior planning approaches that emphasize the generation of safe, effective driving strategies. Sec. 3.3 investigates the interactive relationship between future prediction and behavior planning, highlighting collaborative optimization techniques for complex scenarios. Sec. 4 explores different approaches to data and training paradigms, including supervised and self-supervised learning, and data generation techniques. Sec. 5 examines the application areas and tasks where world models can be applied, discussing the impact of these technologies across diverse domains including perception, prediction, simulation, and system integration. Sec. 6 provides a detailed evaluation of world models for autonomous driving, assessing their effectiveness across various tasks and metrics. Sec. 7 explores open challenges, potential research avenues, and promising directions for further innovation in autonomous driving technologies. Sec. 8 concludes the survey and summarizes key findings, reiterating the importance of robust world models for autonomous driving.
@@ 2 Background
In this section, we first provide a detailed problem formulation for world models in autonomous driving (Sec. 2.1), encompassing two key tasks: future prediction of the physical world and behavior planning for intelligent agents. Then, in Sec. 2.2, we introduce key terminologies and concepts relevant to world models, such as representation spaces, generative models, and spatiotemporal modeling techniques. These aspects lay the foundation for understanding state-of-the-art methods.
@@ 2.1.1 Core Tasks in World Models
The first core task is future prediction of the physical world [29, 27, 17, 30], which involves forecasting the future states of dynamic entities such as vehicles, pedestrians, and traffic elements. This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes. Advanced techniques, such as 4D occupancy prediction and generative models, play a crucial role in addressing this challenge by leveraging multi-modal sensor data and probabilistic forecasting frameworks. The second core task is behavior planning for intelligent agents [18, 31, 19, 32], focusing on generating optimal and feasible trajectories for the ego vehicle. It requires accounting for safety constraints, dynamic obstacles, traffic regulations, and real-time adaptability. Behavior planning is often achieved through a combination of model-based and learning-based approaches that ensure robustness and responsiveness in diverse driving scenarios.
@@ 2.2.2 Generative Models
2.2.3 Spatiotemporal Modeling Techniques
Spatiotemporal modeling addresses how entities evolve in space and time within dynamic driving environments.
Transformers. employ attention mechanisms to capture long-range temporal dependencies and multi-agent interactions, facilitating improved motion prediction and scene understanding [35].
Slot Attention, on the other hand, segments the environment into interpretable entities, enabling more structured representations for downstream tasks such as behavior planning and motion forecasting [11]. These approaches have proven valuable in addressing dynamic and complex driving scenarios. However, developing models that are both computationally efficient and robust in real-time settings remains a core challenge.
This section provides a foundational understanding of world models in autonomous driving, highlighting how representation spaces, generative methods, and advanced modeling techniques together address the critical issues of uncertainty, complexity, and real-time operation.
@@ 3 Taxonomy
In this section, we classify the methodologies presented in the uploaded works into three primary categories, highlighting their unique contributions and interconnections: (1) Future Prediction of the Physical World, (2) Behavior Planning for Intelligent Agents, and (3) Interaction Between Behavior Planning and Future Prediction. This section provides a structured understanding of the diverse approaches in autonomous driving research. Future prediction methods focus on modeling the future states of static structures and dynamic entities, behavior planning optimizes vehicle trajectories, and their integration bridges the gap between environmental understanding and actionable decision-making. These frameworks collectively address the multifaceted challenges of autonomous navigation, driving innovation in safety, efficiency, and robustness.
3.1 Future Prediction of the Physical World
3.1.1 Image Representations
@@
Image representation-based approaches to future prediction in autonomous driving capitalize on advanced generative models to address challenges like limited data availability and complex scene dynamics. Techniques such as diffusion models and generative adversarial networks (GANs) enable the synthesis of high-fidelity images, enriching training datasets and enhancing the robustness of downstream perception and planning modules.
For instance, DrivingDiffusion [14] deploys a diffusion model to generate temporally coherent video frames, creating diverse training scenarios that bolster object detection and motion segmentation in dynamic environments. By employing cascaded generation modules and localized prompts, DrivingDiffusion maintains spatial and temporal consistency, which is pivotal for reliably modeling traffic agents in motion. DriveDreamer [11] further integrates image synthesis and trajectory prediction to simulate realistic urban driving conditions, especially high-density intersections. This focus on interactive elements and unpredictable traffic flows enriches the planning process, enabling vehicles to anticipate and adapt to sudden changes on the road. Meanwhile, ViDAR [17] leverages GANs to synthesize future frames from monocular video sequences, employing temporal consistency layers to mitigate artifacts in moving objects and adapt to varying environmental factors such as weather changes.
Building on these foundations, DrivingWorld [36] proposes a GPT-based video generation framework that unifies detection, prediction, and planning tasks within a single model, thereby streamlining data processing and decision-making. Complementarily, HoloDrive [37] synthesizes 2D–3D multi-modal scenes to reinforce the alignment between image and LiDAR data, enhancing predictive accuracy for more robust reconstruction of complex driving environments. ReconDreamer [38] focuses on the real-time reconstruction of dynamic scenes, emphasizing online restoration techniques that enable responsive modeling of evolving traffic conditions. In simulation environments, Imagine-2-Drive [39] advances high-fidelity world modeling in CARLA, equipping researchers and practitioners with more realistic test scenarios to evaluate autonomous systems under diverse conditions. Finally, Delphi [40] moves toward end-to-end autonomous driving by generating controllable, long-duration videos that facilitate broader generalization in complex traffic scenarios.
@@
Collectively, these image-based generative frameworks significantly strengthen the adaptability and safety of autonomous vehicles. They not only supply synthetic but highly realistic training data for perception tasks, but also refine predictive capabilities and planning strategies, enabling intelligent vehicles to navigate with greater awareness and robustness in rapidly shifting driving contexts.
3.1.2 BEV Representations
@@
A prominent illustration of BEV’s capabilities is GenAD [19], which introduces an end-to-end generative motion prediction pipeline via tokenized scene representations and a Variational Autoencoder (VAE). By capturing interactions between the ego vehicle and surrounding agents in a structured BEV space, GenAD yields accurate trajectory predictions even in cluttered traffic scenarios. Extending this representation to account for evolving semantic occupancy, BEVWorld [27] integrates a diffusion model to predict the geometric and semantic evolution of forthcoming scenes, merging multi-modal inputs (e.g., LiDAR and images) into a continuous latent space. In parallel, UNO [29] tackles 4D occupancy prediction from a purely unsupervised standpoint, leveraging historical signals to generate future point clouds and occupancy grids in a continuous fashion. Its implicit decoder mitigates quantization pitfalls and achieves state-of-the-art outcomes on benchmarks such as Argoverse and nuScenes. Similarly, FIERY [30] exploits spatiotemporal convolutional networks atop BEV inputs to model the dynamic evolution of traffic entities; by integrating probabilistic forecasting, FIERY can output a spectrum of possible future scenarios, which is particularly valuable for safety-critical tasks like instance segmentation and motion prediction.
In addition to these representative works, a host of recent studies further highlights the diversity and adaptability of BEV-based world models. Methods like BevDepth [41] and PowerBEV [42] refine depth estimation and instance prediction within the BEV framework, whereas BevControl [43] and BEVGen [44] push the boundaries of multi-perspective consistency and generative street-view rendering, respectively. At the same time, InfinityDrive [28], GaussianWorld [45], and NeMo [46] expand upon volumetric, generative, and streaming occupancy strategies, illustrating the versatility of BEV-centric approaches when combined with next-generation neural architectures. Collectively, these advances underscore how BEV representations not only offer a unified canvas for perception, prediction, and planning but also significantly improve robustness in dynamic driving environments, paving the way for safer and more reliable autonomous systems.
Generative scene synthesis has recently emerged as a valuable tool for enriching the data landscape used in behavior planning. For instance, BEVGen [44] synthesizes multi-view images from BEV segmentation masks, thereby enhancing data augmentation for tasks such as 3D detection and tracking. By manipulating the bird’s-eye view layout, BEVGen offers a structured way to generate diverse driving scenarios – helping learning-based planners acquire greater robustness and adaptability.
Similarly, BEVControl [43] focuses on simplifying scene editing while maintaining geometric and visual consistency in the top-down view. This tool provides intuitive visual guidance for planning by allowing practitioners to modify complex elements of the scene with minimal effort, thus accelerating the iterative design and testing of learned planning strategies.
Generative scene synthesis also contributes to behavior planning by enriching the data and visual references available. BEVGen [44] synthesizes multi-view images from BEV segmentation masks, enhancing data augmentation for tasks like 3D detection and tracking. BEVControl [43] simplifies scene editing while ensuring geometric and visual consistency, providing intuitive visual guidance for planning. These learning-based approaches underscore the transformative potential of data-driven behavior planning frameworks in addressing the complexities of autonomous driving, enabling safer and more adaptive decision-making in dynamic environments.
3.1.3 OG Representations
@@
Occupancy grid (OG) representations have become a cornerstone for predicting future states in autonomous driving, primarily due to their capacity to encode high-fidelity 3D spatial information. By discretizing the environment into voxel grids, these methods offer a more detailed and geometrically consistent view than projections like BEV, making them especially well-suited for modeling scene evolution and capturing fine-grained interactions over time. However, such high-resolution modeling often incurs substantial memory and computational costs, especially in large-scale or real-time scenarios. The potential for sparse sensor data or irregular updates further complicates prediction accuracy, and controlled scenario generation typically relies on carefully curated trajectory conditions.
A foundational approach in this domain is OccWorld [47], which introduces 4D occupancy grids to jointly predict semantic occupancy and agent trajectories in cluttered driving environments. Building on this groundwork, MUVO [48] integrates multi-modal data (i.e., LiDAR point clouds and camera streams) into high-resolution 3D voxel grids, elevating prediction quality in both vision-based and point cloud tasks. Extending temporal modeling, OccSora [34] leverages diffusion models to generate future occupancy states, enabling controlled scenario simulation for motion-planning pipelines. Further integrating prediction with decision-making, Drive-OccWorld [49] fuses occupancy forecasts with a motion planner, leveraging historical grid states and vehicle trajectories to optimize driving strategies in real-world contexts.
Recent works push OG-based methods toward greater efficiency and scalability. DFIT-OccWorld [50] employs a decoupled dynamic flow mechanism and image-based supervision, achieving end-to-end, non-autoregressive occupancy forecasting while mitigating computational overhead. GaussianWorld [45] reframes OG prediction in a 4D Gaussian space, explicitly modeling scene evolution to enhance streaming performance on datasets like nuScenes without substantial resource demands. Across these advances, learning-based OG techniques consistently prove their worth in motion forecasting, scene reconstruction, and safe route planning. By balancing high-resolution occupancy outputs with emerging strategies for real-time efficiency, the latest research points toward solutions that can seamlessly adapt to the complexity and dynamism of real-world autonomous driving.
3.1.4 PC Representations

### v2

@@ Abstract
Recent breakthroughs in autonomous driving have been propelled by advances in robust world modeling, fundamentally transforming how vehicles interpret dynamic scenes and execute safe decision-making. In particular, world models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics. This paper systematically reviews recent advances in world models for autonomous driving, proposing a three-tiered taxonomy: 1) Generation of Future Physical World, covering image-, BEV-, OG-, and PC-based generation methods that enhance scene evolution modeling through diffusion models and 4D occupancy forecasting; 2) Behavior Planning for Intelligent Agents, combining rule-driven and learning-based paradigms with cost map optimization and reinforcement learning for trajectory generation in complex traffic conditions; 3) Interaction Between Prediction and Planning, achieving multi-agent collaborative decision-making through latent space diffusion and memory-augmented architectures. The study further analyzes training paradigms including self-supervised learning, multimodal pretraining, and generative data augmentation, while evaluating world models’ performance in scene understanding and motion prediction tasks. Future research must address key challenges in self-supervised representation learning, long-tail scenario generation, and multimodal fusion to advance the practical deployment of world models in complex urban environments. Overall, our comprehensive analysis provides a theoretical framework and technical roadmap for harnessing the transformative potential of world models in advancing safe and reliable autonomous driving solutions.
@@ 1.1 Overview
Within this context, constructing robust and stable world models [12] has emerged as a foundational element. The notion of a world model involves creating a high-fidelity representation of the driving environment – encompassing static structures (e.g., roads, buildings) and dynamic entities (e.g., vehicles, pedestrians) [3, 8]. A comprehensive world model continuously captures semantic and geometric information while updating these representations in real-time, thereby informing downstream tasks such as physical world prediction [13, 14]. Recent advances integrate multi-sensor data to refine these representations, such as generative approaches [15, 16] that simulate the physical world for training that unify heterogeneous sensor inputs into consistent top-down perspectives [17, 18].
@@ 1.2 Contributions
Guided by the principle that the world model is central to the understanding of dynamic scenes, this survey aims to provide a comprehensive, structured review of existing methodologies. We categorize state-of-the-art research into three key areas: Generation of Future Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [11, 26]; Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [13, 14]; Interaction Between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [27, 28, 19]. Specifically, we provide:
  • An In-Depth Analysis of Future Prediction Models: We discuss how Image-/BEV-/OG-/PC-based generation methods achieve geometric and semantic fidelity in dynamic scenes, including 4D occupancy forecasting and diffusion-based generation.
@@
§1§2§3§4§5§6§7§8§2§3.1§3.2§3.3§4§5
Figure 1: Structure of the overall review. The top row outlines the paper’s organization. The second and third rows illustrate the background and key components – generation of future physical world, behavior planning for intelligent agents, and the interaction between them. The fourth row highlights various methodologies for training models in autonomous driving, covering self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation. The bottom row showcases four application areas for world models in autonomous driving: scene understanding, motion prediction, simulation, and end-to-end driving.
@@ 1.3 Structure
A summary of the structure of this paper can be found in Fig. 1, which is presented as follows: Sec. 1 introduces the significance of world models in autonomous driving and outlines the societal and technical challenges they address. Sec. 2 provides a comprehensive background on the formulation and core tasks of world models in autonomous driving, specifically focusing on the future prediction of the physical world and behavior planning for intelligent agents. Sec. 3 details the taxonomy of methods: Sec. 3.1 delves into methods for generation of future physical world, discussing physical world evolution of dynamic objects and static entities. Sec. 3.2 discusses advanced behavior planning approaches that emphasize the generation of safe, effective driving strategies. Sec. 3.3 investigates the interactive relationship between future prediction and behavior planning, highlighting collaborative optimization techniques for complex scenarios. Sec. 4 explores different approaches to data and training paradigms, including supervised and self-supervised learning, and data generation techniques. Sec. 5 examines the application areas and tasks where world models can be applied, discussing the impact of these technologies across diverse domains including perception, prediction, simulation, and system integration. Sec. 6 provides a detailed evaluation of world models for autonomous driving, assessing their effectiveness across various tasks and metrics. Sec. 7 explores open challenges, potential research avenues, and promising directions for further innovation in autonomous driving technologies. Sec. 8 concludes the survey and summarizes key findings, reiterating the importance of robust world models for autonomous driving.
@@ 2 Background
In this section, we first provide a detailed problem formulation for world models in autonomous driving (Sec. 2.1), encompassing two key tasks: generation of future physical world and behavior planning for intelligent agents. Then, in Sec. 2.2, we introduce key terminologies and concepts relevant to world models, such as representation spaces, generative models, and spatiotemporal modeling techniques. These aspects lay the foundation for understanding state-of-the-art methods.
@@ 2.1.1 Core Tasks in World Models
The first core task is generation of future physical world [30, 28, 18, 31], which involves forecasting the future states of dynamic entities such as vehicles, pedestrians, and traffic elements. This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes. Advanced techniques, such as 4D occupancy prediction and generative models, play a crucial role in addressing this challenge by leveraging multi-modal sensor data and probabilistic forecasting frameworks. The second core task is behavior planning for intelligent agents [19, 32, 20, 33], focusing on generating optimal and feasible trajectories for the ego vehicle. It requires accounting for safety constraints, dynamic obstacles, traffic regulations, and real-time adaptability. Behavior planning is often achieved through a combination of model-based and learning-based approaches that ensure robustness and responsiveness in diverse driving scenarios.
@@ 3 Taxonomy
In this section, we classify the methodologies presented in the uploaded works into three primary categories, highlighting their unique contributions and interconnections: (1) Generation of Future Physical World, (2) Behavior Planning for Intelligent Agents, and (3) Interaction Between Behavior Planning and Future Prediction. This section provides a structured understanding of the diverse approaches in autonomous driving research. Future prediction methods focus on modeling the future states of static structures and dynamic entities, behavior planning optimizes vehicle trajectories, and their integration bridges the gap between environmental understanding and actionable decision-making. These frameworks collectively address the multifaceted challenges of autonomous navigation, driving innovation in safety, efficiency, and robustness.
3.1 Generation of Future Physical World
3.1.1 Image-based Generation
@@
Image representation-based approaches to future prediction in autonomous driving draw on advanced generative models, such as diffusion models, to address issues like limited labeled data and highly dynamic environments. By synthesizing high-fidelity images, these methods expand training datasets and bolster the resilience of downstream perception and planning modules.
DriveDreamer [11] uses real-world driving data for controllable video generation in high-density urban traffic; DriveDreamer-2 [36] incorporates LLM-based prompts for customized multi-view video generation; DriveDreamer4D [26] leverages world model priors to produce spatial-temporally consistent 4D driving videos; ReconDreamer [37] integrates online restoration for accurate video reconstruction of dynamic scenes; WorldDreamer [38] predicts masked tokens to enable text-to-video and action-to-video generation; CarDreamer [39] provides an open-source platform supporting multi-modal video generation for autonomous driving.
DrivingDiffusion [15] employs a latent diffusion model for layout-guided, multi-view driving scene video generation; Delphi [40] applies a controllable, long-duration diffusion-based approach to boost planning in autonomous driving videos. DrivingWorld [41] introduces a GPT-based framework for extended, high-fidelity driving video generation; GAIA-1 [16] uses unsupervised sequence modeling that integrates video, text, and action inputs for contextual driving video generation; DRIVESIM [42] leverages multimodal large language models to achieve robust, causal driving video predictions. HoloDrive [43] fuses camera and LiDAR data for future-frame driving video generation; BevWorld [28] applies cross-modal BEV learning for dynamic scene video generation; BEVGen [44] synthesizes street-level videos from semantic BEV layouts; Drive-WM[45] unifies visual forecasting and planning for consistent driving video generation. DriveArena [46] creates a closed-loop simulation for generative driving video; DrivingSphere [47] builds a 4D dynamic environment for multi-view driving video; Imagine-2-Drive [48] merges a high-fidelity world model with a diffusion-based policy network for diverse trajectory video generation; Vista [49] programmatically renders complex driving scenarios as high-quality videos; SimGen [50] conditions on simulators to produce varied driving videos for real-world generalization.LAW [51] trains a latent world model to predict future scene videos for end-to-end driving; Popov et al. [33] propose incremental latent data generation to combat environmental shifts in driving videos.
@@
Overall, these image-centric generative frameworks significantly enhance the adaptability and safety of autonomous vehicles. Beyond providing synthetic yet exceptionally realistic data for perception tasks, they also refine predictive capabilities and planning strategies, enabling intelligent vehicles to navigate rapidly changing road conditions with heightened awareness and robustness.
3.1.2 BEV-based Generation
@@
CarFormer [52] incorporates an object-centric BEV representation, splitting vehicles, pedestrians, and other entities into independent slots for more precise trajectory analysis. Building on similar structured representations, MILE [32] integrates a world model into imitation learning for multi-modal BEV predictions, while Popov et al. [33] further extend this pipeline by introducing latent-space generative models to mitigate covariate shift, ensuring semantic consistency between BEV and perspective views for enhanced robustness. GenAD [20] reframes autonomous driving as a generative modeling process, encoding scene elements in a BEV domain and sampling future trajectories through structured latent spaces. In parallel, UNO [30] focuses on unsupervised 4D occupancy field learning in BEV, enabling seamless adaptation to point cloud and BEV semantic occupancy forecasting, whereas FIERY [31] leverages spatiotemporal convolutions atop BEV inputs to predict multi-modal future instance trajectories under probabilistic uncertainty. PowerBEV [53] targets efficiency and stability by employing parallel multi-scale modules and flow-guided post-processing to reduce redundancy in BEV forecasting, achieving state-of-the-art performance on benchmarks like NuScenes with minimal model overhead. In contrast, BEVControl [54] addresses multi-perspective consistency by editing a BEV sketch layout, ensuring coherent geometry from top-down to perspective views—thereby offering an intuitive, flexible tool for scene manipulation and planning validation without directly outputting future trajectories.
Aside from the aforementioned methods, to achieve substantial improvement in BEV semantic occupancy forecasting, there is another line of research that attempts to introduce a pre-text task called visual point cloud forecasting. ViDAR [18] proposes a novel pre-training framework for autonomous driving through a task termed visual point cloud forecasting—predicting future 3D point cloud sequences from historical visual inputs. With the aid of ViDAR, performance improvements are achieved in BEV generation.
3.1.3 OG-based Generation
@@
Occupancy grid (OG) representations have become a cornerstone for predicting future states in autonomous driving, primarily due to their capacity to encode high-fidelity 3D spatial information. By discretizing the environment into voxel grids, these methods offer a more detailed and geometrically consistent view than projections like BEV, making them especially well-suited for modeling scene evolution and capturing fine-grained interactions over time. However, such high-resolution modeling often incurs substantial memory and computational costs, especially in large-scale or real-time scenarios.
OG-based Generation (i.e., occupancy forecasting) originates from forecasting semantic occupancy grids on bird’s-eye view (BEV) [55, 56, 57, 58, 59, 60]. It is employed to predict how the surrounding occupancy will evolve in the near future beyond the current moment. Then gradually expand to predict 4D occupancy grids. For instance, Occ4cast [61] aims to forecast a sequence of dense and completed occupancy grids under the Eulerian specification, offering a more comprehensive perception in complex dynamic environments. Recent OG-based Generation approaches begin to closely align with world models, showing a trend toward integrating multi-view images, language, and actions to achieve advanced 4D occupancy forecasting and planning. They adopt multi-modal sensors, generative architectures, and spatiotemporal tokenization strategies to comprehensively model future occupancy for decision-making in autonomous driving. Specifically, MUVO [62] pioneers multi-modal sensor fusion (camera + LiDAR) for actionable 3D occupancy prediction, while Cam4DOcc [63] establishes the first camera-only 4D occupancy benchmark with standardized protocols. Building on these, OccWorld [64] introduces GPT-style autoregressive generation for scene token evolution, and DFIT-OccWorld [65] streamlines training via decoupled dynamic-static voxel warping. Concurrently, DOME [66] adopts diffusion transformers for controllable long-horizon forecasting, whereas DriveWorld [19] and Drive-OccWorld [67] integrate memory-augmented world models with BEV-based planning, bridging occupancy forecasting and trajectory optimization. GaussianWorld [68] reformulates scene dynamics in 3D Gaussian space, enabling ego-motion-aligned prediction, while OccLLaMA [69] unifies vision-language-action modalities via occupancy tokenization. OccSora [35] extends this with diffusion-based 4D generation conditioned on trajectories, and RenderWorld [70] achieves vision-only efficiency via Gaussian Splatting and disentangled VAE encoding, culminating in a cost-effective, end-to-end autonomy pipeline.
3.1.4 PC-based Generation